Jenelle bought a home for $470,000, paying 20% as a down payment, and financing the rest at 6.8% interest for 30 years. Round your answers to the nearest cent.
How much money did Jenelle pay as a down payment? $
What was the original amount financed? $
What is her monthly payment? $
If Jenelle makes these payments every month for thirty years, determine the total amount of money she will spend on this home. Include the down payment in your answer. $
Answer by Boreal(15235) 
 down payment is 20% of $470,000 or $94,000, so P=the difference or $376,000 originally financed
Monthly payment is P*r/n(1+(r/n))^nt)/(1+(r/n)^nt -1); r=0.068, n=12, nt=360, round at end
370,000(0.068)*7.646/6.646=$2412.12
Total cost is: $94,000 down payment
+2421.12*360=$962,363.20
| |
|